OpenStudy (anonymous):

1-(cos^2x/1+sinx) write as a constant, a single trigonometric funciton, or a power of a trigonometric function.

OpenStudy (anonymous):

1 - (cos^2 x)/(1 + sin x) =[ (1 + sin x) - (1 - sin^2 x)](1 + sin x) = = (sin x + sin^2 x)/(1 + sin x) = [sin x*(1 + sin x)]/(1 + sin x = sin x.